The weights of certain machine componets are normally distrubited with a mean of 8.75g and a standard deviation of 0.07g. Find the two weights that seperate the top 3% and the bottom 3%. These weights could serve as limit used to identify which componets should be rejected.
Answer by stanbon(75887)

The weights of certain machine componets are normally distrubited with a mean of 8.75g and a standard deviation of 0.07g. Find the two weights that seperate the top 3% and the bottom 3%. These weights could serve as limit used to identify which componets should be rejected.
--------------------------------
Use your z-chart or your calculator to find the "z" value that separates the top 3% and the bottom 3%.
That value is z= 1.88 and z=-1.88
-------------------------------
Then find the weights that correspond to those z values:
Let the weight be "x":
z = (x-mu)/sigma
1.88 = (x-8.75)/0.07
x= 0.07*1.88+8.75
x = 8.8816 lbs is the weight that separates the top 3%.
---------------
-1.88 = (x-mu)/sigma
x = 0.07(-1.88)+8.75
x = 8.6184 lbs is the weight that separates the bottom 3%
==================
Cheers,
Stan H.